OpenStudy (anonymous):

Determine the slope from the data in the table below. x -2 , -1, 0, 1 y 2, 1.5, 1, 0.5 Select one: a. 1/2 b. -1/2 c. 2 d. -2

OpenStudy (texaschic101):

(-2,2) x1 = -2 and y1 = 2 (-1,1.5) x2 = -1 and y2 = 1.5 slope formula : (y2 - y1) / (x2 - x1) sub your points in and what do you get ?

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

its B right?

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) slope = (1.5 - 2) / (-1 - (-2) slope = -.5 / (-1 + 2) slope = -0.5/1 or just -0.5 which is the same as -1/2.....it is B

OpenStudy (anonymous):

cool thanks.

OpenStudy (texaschic101):

no problem :)